OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 10 × 5 + 3 ÷ 3 - (4 + 3)

pooja195 (pooja195):

Parathese Exponents Multiplacation Divison Addition Subtraction

OpenStudy (anonymous):

???????????

pooja195 (pooja195):

Those are the order of operations :) Ok so paratheses 4+3=?

OpenStudy (anonymous):

7

pooja195 (pooja195):

next multiply 10 x 5 =?

OpenStudy (anonymous):

50

pooja195 (pooja195):

division 3/3 = ?

OpenStudy (anonymous):

1

pooja195 (pooja195):

addtion 50 + 7 + 1 =?

OpenStudy (anonymous):

58

OpenStudy (anonymous):

what about the subtraction?

pooja195 (pooja195):

wait let me set this up correctly

OpenStudy (anonymous):

ok

pooja195 (pooja195):

\[(10\times5)+(3/3)-(4+3)\]

OpenStudy (anonymous):

7-1=6 6+50=56?

pooja195 (pooja195):

where did u get 6?

OpenStudy (anonymous):

4+3=7 then minus 1

pooja195 (pooja195):

start from th beggening 10 x 5= 50 + 1=51-7=?

OpenStudy (anonymous):

44

pooja195 (pooja195):

yep thats your answer :)

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

your welcome :) also welcome to openstudy!!! :D